# WHAT-IF: Exploring Branching Narratives by Meta-Prompting Large Language Models

### v1

@@ Abstract
WHAT-IF—Writing a Hero’s Alternate Timeline through Interactive Fiction—is a system that uses zero-shot meta-prompting to create branching narratives from a prewritten story. Played as an interactive fiction (IF) game, WHAT-IF lets the player choose between decisions that the large language model (LLM) GPT-4 generates as possible branches in the story. Starting with an existing linear plot as input, a branch is created at each key decision taken by the main character. By meta-prompting the LLM to consider the major plot points from the story, the system produces coherent and well-structured alternate storylines. WHAT-IF stores the branching plot tree in a graph which helps it to both keep track of the story for prompting and maintain the structure for the final IF system. A video demo of our system can be found here: https://youtu.be/8vBqjqtupcc.
@@ 1 Introduction
In this paper, we present WHAT-IF (Writing a Hero’s Alternate Timeline through Interactive Fiction)111https://what-if-game.github.io/, a system for controllable, branching narrative generation in English. Starting with a human-written plot, WHAT-IF produces distinct changes in the plot by considering alternate actions the main character could have taken. We create these choice points by having a large language model (LLM) generate a new sub-goal for the character, which significantly changes subsequent plot points in the story. We use this LLM-driven branching narrative to create an interactive fiction system (Fig. 1) allowing the user to choose which branches to follow in the story tree.
In contrast with previous LLM-driven interactive fiction systems like AI Dungeon222https://aidungeon.com/, which rely entirely on the user to move the story along, WHAT-IF generates more constrained stories where the user has a degree of control but the stories generated by the LLM maintain a high degree of thematic consistency with the original story, even with many possible endings.
@@
Although we present the system as an interactive fiction game, we believe that WHAT-IF could be used to help authors explore alternative paths in storylines they are writing. See Swanson et al. (2021); Yuan et al. (2022); Zhu et al. (2023a); Gero et al. (2023) for some examples of writing aids using LLMs. Park et al. (2023) even look at using ChatGPT to help author IFs.
In the rest of this paper, we will discuss related work about using LLMs for interactive fiction and story generation, describe the architecture of WHAT-IF, and then give example narratives that were generated.
2 Related Work: Storytelling with LLMs
@@
On the other hand, large language models (LLMs) yield flexible, grammatical natural language generation. While the previous generation of LLMs (GPT-2 era) had mixed results with story generation See et al. (2019), modern LLMs are telling much better stories and are almost generating at a human level Xie et al. (2023).
There has been plenty of work on controlled story generation with LLMs Lal et al. (2021); Razumovskaia et al. (2024); Yang et al. (2022); Qin and Zhao (2022); Zhu et al. (2023b). Most of this work consists of higher-level “planning” with LLMs, where an outline or set of keywords is expanded into a longer-form story. There have also been recent works using retrieval-augmented LLMs for story generation Zhu et al. (2023a); Wen et al. (2023). However, people have only just begun to explore mixing more traditional AI planning techniques with LLMs, namely by generating components that would be fed into a planner or using pre-conditions and effects in a similar manner to a planner Martin (2021); Simon and Muise (2022); Kelly et al. (2023); Ye et al. (2023).
LLMs have also been used for story-generation–related tasks, including generating story worlds Johnson-Bey et al. (2023) and doing story sifting Méndez and Gervás (2023)—that is, creating a story from a collection of facts.
@@ 2.1 Interactive Narrative Generation
In line with story generation’s planning history, there have been plenty of interactive narrative systems using planning methods Riedl and Bulitko (2013); Sanghrajka (2019); Ware and Siler (2021). So far, neural IF work has included tasks such as generating interactive fiction worlds Ammanabrolu et al. (2020), scenes Kumaran et al. (2023), branching quests De Lima et al. (2021), or choices for the user Harmon and Rutman (2023). There has been a collection of work looking at interactive fiction through roleplaying as well, using recurrent language models Louis and Sutton (2018), reinforcement learning Martin et al. (2018); Ammanabrolu et al. (2021), and LLMs Callison-Burch et al. (2022); Ashby et al. (2023); Wang et al. (2024); Shao et al. (2023); Yan et al. (2023).
There are a couple of systems we’d like to point out in particular to compare to our system WHAT-IF. Pei et al. (2024)’s SWAG story generation system uses an LLM to both generate actions (story events) and then select the best action using Direct Preference Optimization (DPO) Rafailov et al. (2023). Although we are not selecting the action—the user is, we are relying on the LLM to generate story branches. However, instead of finetuning on pre-written stories, WHAT-IF uses zero-shot prompting to generate new events.
Similarly, like Harmon and Rutman (2023)’s works, we use LLMs to generate branching narratives by exploring character choices. Our system expands on this work by, instead of using story-specific questions as part of the prompt, WHAT-IF uses meta-prompts so that the system can ask itself story-specific questions to aid in generation.
@@ 3.1 Initializing the Branching Plot Tree
The tree is constructed as follows (illustrated in Figure 4):
Each node contains the current state () and goal () of the character, the key decision () taken in the original storyline, and the alternate decision () given the same state and goal that would lead to an alternate storyline.
Each edge contains a list of three sentences corresponding to events events that lead the character to move from one node to another: (1) the decision made by the character— or as full sentences , (2) the event resulting from the decision & leading to the next state, and (3) the next state of the character . We will refer to these as edge events.
@@
In this phase (#1 & #5 in Figure 3), we prompt GPT-4 to summarize the story into a storyline in the branching plot tree and imagine the alternate decision at each node. A storyline is defined as a linear traversal of the tree from the root node to one of the leaf nodes. The strong causal relationship among these variables ensures it can capture meaningful decisions, sometimes hidden, in a plot that propels the story forward. Note that extracting events from all edges of a storyline would give us the full plot summary.
@@ 3.2 Extracting Key Events
A well-known story structure is the Three-act Structure that divides the story into three phases: Setup, Confrontation, and Resolution Field (1979). The backbone of this structure is three major plot points: Inciting Incident, Crisis, and Climax. These major plot points force the protagonist to experience and grow and drive the tension up and down, leading to an engaging story.
Therefore, we prompt GPT-4 to identify three key events corresponding to the three major plot points (#2 in Figure 3) from all of the edge events in a storyline and save them for later use in the prompt generation.
@@ 3.3 Generating Meta-Prompts
Although an ideal story should show and not tell, we observed that the LLM is prone to following instructions too explicitly. For example, the prompt of “the alternate storyline should contain new challenges…” would lead to events such as “Tony faces a new challenge of…”. Thus, we adopted a meta-prompting approach (i.e., prompting to generate prompts) which has shown to be an effective method to elicit better responses from LLMs de Wynter et al. (2023); Suzgun and Kalai (2024).
@@ 3.4 Branching & Merging the Tree
Generate a list of new events that continue from the previous (extracted) events. (#4 in Figure 3)
@@
Create a branch from the list of new events (#5).
@@
Merge the branches of the previous events (the old plot tree) and new events to form a new branching plot tree (#6).
@@ 3.5 Narrating the Game
After our binary branching plot tree is complete, we have possible story endings in our game, with being the number of nodes/events in the original storyline. WHAT-IF traverses the tree to generate the narration for each node by prompting GPT-4 to narrate all events of the edge pointing towards the node from the previous time step. After that, it would narrate the state and goal of the character. Lastly, it provides two decisions and for the player to choose from. After the player makes a decision, WHAT-IF continues narration using the corresponding node in the tree. Lastly, we organize all narration into an Ink file—a scripting language developed by Inkle333https://www.inklestudios.com/ for writing interactive narratives—and use InkJS444https://github.com/y-lohse/inkjs to compile and generate the game online.
@@ 4 Qualitative Analysis & Discussion
To evaluate the effectiveness of our system, we use the plot of Iron Man taken from IMDB as the original story and compare the alternate storylines generated by GPT-4 using vanilla prompts and WHAT-IF prompts. Examples of the vanilla and WHAT-IF prompts and their corresponding output can be found in Appendix B.
@@ 4.1 Improving Branch Choices
(1) Non-choices. Since vanilla prompting considers every event as a potential branching point, any events that don’t involve decisions from the main character produce illogical branching events such as “Stark decides to get ambushed by the Ten Rings”. Although this aligns with the original event “Stark is critically wounded and captured by terrorists who demand he build a Jericho missile for them.”, this is not a decision Tony can make.
@@
(3) Retreat decisions. In interactive fiction writing, there is often a retreat option where the player decides to stay out of a situation. However, this would not be an ideal branching point that is meaningful to the story as it does not move the story forward. For example, “Tony decides not to save the villagers.

### v3

@@ Abstract
WHAT-IF—Writing a Hero’s Alternate Timeline through Interactive Fiction—is a system that uses zero-shot meta-prompting to create branching narratives from a prewritten story. Played as an interactive fiction (IF) game, what-if lets the player choose between decisions that the large language model (LLM) GPT-4 generates as possible branches in the story. Starting with an existing linear plot as input, a branch is created at each key decision taken by the main character. By meta-prompting the LLM to consider the major plot points from the story, the system produces coherent and well-structured alternate storylines. what-if stores the branching plot tree in a graph which helps it to both keep track of the story for prompting and maintain the structure for the final IF system. A demo of what-if can be found at https://what-if-game.github.io/.
@@ 1 Introduction
In this paper, we present WHAT-IF (Writing a Hero’s Alternate Timeline through Interactive Fiction), a system for controllable, branching narrative generation in English. Starting with a human-written plot, what-if produces distinct changes in the plot by considering alternate actions the main character could have taken. We create these choice points by having a large language model (LLM) generate a new sub-goal for the character, which significantly changes subsequent plot points in the story. We use this LLM-driven branching narrative to create an interactive fiction system (Fig. 1) allowing the user to choose which branches to follow in the story tree.
In contrast with previous LLM-driven interactive fiction systems like AI Dungeon Walton (2019) which rely entirely on the user to move the story along, what-if generates more constrained stories. The user has a degree of control, but the stories generated by the LLM maintain a high degree of thematic consistency with the original story, even with many possible endings.
@@
Although we present the system as an interactive fiction game, we believe that what-if has potential applications for helping authors explore alternative paths in storylines they are writing. For example, LLMs have been used as writing aids Swanson et al. (2021); Yuan et al. (2022); Zhu et al. (2023a); Gero et al. (2023), and have even been shown to help people author IFs Park et al. (2023).
In the rest of this paper, we will discuss related work about using LLMs for interactive fiction and story generation, describe the architecture of what-if, and then give example narratives that were generated. The code for generating branching narratives with what-if can be found at https://github.com/what-if-game/what-if-game.github.io.
2 Related Work
2.1 Storytelling with LLMs
@@
On the other hand, LLMs yield flexible, grammatical natural language generation. While early LLMs had mixed results with story generation See et al. (2019), more modern LLMs are telling better stories and are able to generate short stories at a human level Xie et al. (2023); Wang et al. (2025).
There has been plenty of work on controlled story generation with LLMs Lal et al. (2021); Razumovskaia et al. (2024); Yang et al. (2022); Qin and Zhao (2022); Zhu et al. (2023b); Kumaran et al. (2024); Wang et al. (2025); Xia et al. (2025); Zimmerman et al. (2025). Most of this work consists of higher-level “planning” with LLMs, where an outline or set of keywords is expanded into a longer-form story. There have also been recent works using retrieval-augmented LLMs for story generation Zhu et al. (2023a); Wen et al. (2023). However, people have only just begun to explore mixing more traditional AI planning techniques with LLMs, namely by generating components that would be fed into a planner or using pre-conditions and effects in a similar manner to a planner Martin (2021); Simon and Muise (2022); Kelly et al. (2023); Ye et al. (2023); Farrell and Ware (2024).
@@ 2.2 Interactive Narrative Generation
There have been interactive narrative systems using planning methods as well Riedl and Bulitko (2013); Sanghrajka (2019); Ware and Siler (2021). So far, neural IF work has included tasks such as generating interactive fiction worlds Ammanabrolu et al. (2020); Zhou et al. (2025), scenes Kumaran et al. (2023), branching quests De Lima et al. (2021), or choices for the user Harmon and Rutman (2023). We take inspiration from De Lima et al. (2021) and adapt their quest tree planning methods to generating a plot tree with an LLM.
There has been a collection of work looking at interactive fiction through roleplaying as well, using recurrent language models Louis and Sutton (2018), reinforcement learning Martin et al. (2018); Ammanabrolu et al. (2021), and LLMs Callison-Burch et al. (2022); Ashby et al. (2023); Wang et al. (2024); Shao et al. (2023); Yan et al. (2023).
There are several systems that are of note as comparisons to our system what-if. Pei et al. (2024)’s SWAG story generation system uses an LLM to both generate actions (story events) and then select the best action using Direct Preference Optimization (DPO) Rafailov et al. (2023). Although we are not selecting the action—the user is, we are relying on the LLM to generate story branches. However, instead of finetuning on pre-written stories, what-if uses zero-shot prompting to generate new events.
Similarly, like Harmon and Rutman (2023)’s works, we use LLMs to generate branching narratives by exploring character choices. However, instead of using story-specific questions as part of the prompt, what-if uses meta-prompts so that the model can ask itself story-specific questions to aid in generation.
@@ 3 System Architecture
For each component of the pipeline, we use GPT-4, which was the state-of-the-art at the time of creating what-if. Each instance of the model was used out-of-the-box (not finetuned) to test how well it could be pushed to generated well-formed alternative storylines for stories and characters it has most likely memorized. This is also the reasoning behind choosing well-known stories.
@@ 3.1 Initializing the Branching Plot Tree
The tree (illustrated in Figure 4) is constructed as follows:
Each node contains the current state () and goal () of the character, the key decision () taken in the original storyline, and an alternate decision () which would lead to an alternate storyline given the same and .
Each edge contains a list of three sentences corresponding to events that lead the character to move from one node to another: (1) the decision made by the character— or as full sentences, (2) the event resulting from the decision & leading to the next state, and (3) the next state of the character . We will refer to these as edge events.
@@
In this phase ( & in Figure 3), we prompt GPT-4 to summarize the story into a storyline in the branching plot tree and imagine the alternate decision at each node. A storyline is defined as a linear traversal of the tree from the root node to one of the leaf nodes. The strong causal relationship among these variables ensures it can capture meaningful decisions, sometimes hidden, in a plot that propels the story forward. Note that extracting events from all edges of a storyline would give us the full plot summary.
@@ 3.2 Extracting Key Events
A well-known story structure is the Three-Act Structure Field (1979) that divides the story into three phases: Setup, Confrontation, and Resolution. The backbone of this structure is three major plot points: Inciting Incident, Crisis, and Climax. These major plot points force the protagonist to experience and grow and drive the tension up and down, leading to an engaging story.
@@ 3.3 Generating Meta-Prompts
Although an ideal story should show and not tell, we observed that the LLM is prone to following instructions too explicitly. For example, the prompt “the alternate storyline should contain new challenges…” would lead to events such as “Tony faces a new challenge of…”. Thus, we adopted a meta-prompting approach (i.e., prompting to generate prompts) which has shown to be an effective method to elicit better responses from LLMs de Wynter et al. (2023); Suzgun and Kalai (2024).
@@ 3.4 Branching & Merging the Tree
Generate a list of new events that continue from the previous (extracted) events. ( in Figure 3)
@@
Create a branch from the list of new events. ( )
@@
Merge the branches of the previous events (the old plot tree) and new events to form a new branching plot tree. ( )
@@ 3.5 Narrating the Game
After our binary branching plot tree is complete, we have possible story endings in our game, with being the number of nodes/events in the original storyline. what-if traverses the tree to generate the narration for each node by prompting GPT-4 to narrate all events of the edge pointing towards the node from the previous time step. After that, it would narrate the state and goal of the character. Lastly, it provides two decisions and for the player to choose from. An example of this generation for a single node can be found in Table 1. A full example can be found in Appendix A.
After the player makes a decision, what-if continues narration using the corresponding node in the tree. Lastly, we organize all narration into an Ink file—a scripting language developed by Inkle111https://www.inklestudios.com/ for writing interactive narratives—and use InkJS222https://github.com/y-lohse/inkjs to compile and generate the game online.
@@ 4 Qualitative Analysis & Discussion
During the design of the system, we used the plot of Iron Man taken from IMDB as the original story and compared the alternate storylines generated by GPT-4 using vanilla prompts and what-if prompts. Examples of the vanilla and what-if prompts and their corresponding output can be found in Appendix B. We share some of our insights of this comparison below.
@@ 4.1 Improving Branch Choices
(1) Non-choices. Since vanilla prompting considers every event as a potential branching point, any events that do not involve decisions from the main character produce illogical branching events such as “Stark decides to get ambushed by the Ten Rings”. Although this aligns with the original event “Stark is critically wounded and captured by terrorists who demand he build a Jericho missile for them.”, this is not a decision Tony can make.
@@
(3) Retreat decisions. In interactive fiction writing, there is often a retreat option where the player decides to stay out of a situation. However, this would not be an ideal branching point that is meaningful to the story as it does not move the story forward. For example, “Tony decides not to save the villagers.”